Recreational marijuana was legalized in the state of Washington in July of 2014 creating a new source of revenue for the government. July 1, 2015, the state government raised excise taxes paid by marijuana retailers on final sales leading to the following effects on prices and quantity sold:

Date

Quantity Sold (average grams per day)

Price (per gram)

June 2015 (pre tax change)

310.0

13.18

July 2015

307.0

13.48

August 2015

303.1

13.48

The above example is adapted from a recent study into the effects of an increase in taxes on the marijuana market in Washington (Hansen, Miller, & Weber, 2017). The authors can estimate the price elasticity of demand from these data points because the sudden tax change created a ceteris paribus situation, where all other supply and demand conditions were held fixed, similar to the way our models assume in the textbook.

In that spirit, interpret the excise tax increases as an increase in input costs for marijuana dispensaries holding all other supply and demand characteristics fixed.

Comparing June 2015 to July 2015, using the arc formula for price elasticity of demand: What is the price elasticity of demand? If the business didn’t have to pay the sales tax to the government, what affect would the price increase have had on their revenue? Type your response in the box below.
Answer #1

The "Arc Formula" for Price Elasticity of Demand is given as,

Price Elasticity of Demand = [ (Q2 - Q1) / Midpoint Q ] / [ ( P2 - P1) / Midpoint P ]

Q1 = 310 ; Q2 = 307; P1 = 13.18 ; P2 = 13.48

Midpoint Q = (Q1 + Q2) / 2 = (310 + 307) / 2 = 308.5

Midpoint P = (P1 + P2) / 2 = (13.18 + 13.48) / 2 = 13.33

Price Elasticity of Demand = [ (307 - 310) / 308.5 ] / [ (13.48 - 13.18) / 13.33 ]

Price Elasticity of Demand = -0.432

Since the price elasticity of demand is less than 1, demand for marijuana is inelastic.

Note:

When demand is inelastic, price and total revenue move in the same direction.

In other words, if businesses didn't have to pay the sales tax to the government, the price increase would have translated into higher revenues.